Major League Apparel has two classes of stock authorized: 5%, $10 par preferred, and $1 par value common. The following transactions affect stockholders’ equity during 2021, its first year of operations:

January 2 Issue 100,000 shares of common stock for $69 per share.
February 14 Issue 59,000 shares of preferred stock for $11 per share.
May 8 Purchase 10,000 shares of its own common stock for $59 per share.
May 31 Resell 5,000 shares of treasury stock for $64 per share.
December 1 Declare a cash dividend on its common stock of $0.30 per share and a $29,500 (5% of par value) cash dividend on its preferred stock payable to all stockholders of record on December 15. The dividend is payable on December 30. (Hint: Dividends are not paid on treasury stock.)
December 30 Pay the cash dividends declared on December 1.

2. Prepare the stockholders’ equity section of the balance sheet as of December 31, 2021. Net income for the year was $489,000. (Amounts to be deducted should be indicated by a minus sign.)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Major League Apparel
Workings
Common Stock Number Note
Shares Issued on 2nd Jan       100,000.00 A
Sale price                 69.00 B
Par Value per share                   1.00 C
Paid in capital in excess of par- Common Stock                 68.00 D=B-C
Sale value 6,900,000.00 E=B*A
Par Value       100,000.00 F=C*A
Paid in capital in excess of par- Common Stock 6,800,000.00 G=D*A
Common Stock- Dividend
Shares Issued on 2nd Jan       100,000.00 See A
Less: Reacquired on May 8         10,000.00
Add: Sold reacquired shares on May 31            5,000.00
Number of shares on which dividend paid         95,000.00 H
Dividend per share                   1.00 I
Common Stock- Dividend         95,000.00 J=H*I
Treasury Stock Number
Number of shares reacquired on May 8         10,000.00 K
Price paid per share                 59.00 L
Total amount paid       590,000.00 M=K*L
Less: Reissued shares on May 31            5,000.00 N
Number of shares as Treasury Stock as on Dec 31           5,000.00 O=K-N
Price paid per share                 59.00 See L
Value of Treasury Stock as on Dec 31       295,000.00 P=O*L
Paid in capital in excess of par- Common Stock
From issue of shares on 2nd Jan    6,800,000.00 See G
Reissued 5,000 shares @ $ 64 on May 31       320,000.00 Q
Less: Price paid for 5,000 shares @ $ 59 on May 8       295,000.00 R
Paid in capital in excess of par- Common Stock         25,000.00 S=Q-R
Closing Paid in capital in excess of par- Common Stock 6,825,000.00 O=G+S
Preferred Stock Number Note
Shares Issued on Feb 2         59,000.00 T
Sale price                 11.00 U
Par Value per share                 10.00 V
Paid in capital in excess of par- Preferred Stock                   1.00 W=U-V
Sale value       649,000.00 X=T*U
Par Value       590,000.00 Y=T*V
Paid in capital in excess of par- Preferred Stock         59,000.00 Z=T*W
Retained Earnings
Income Summary       489,000.00
Less:
Preferred Stock- Dividend         29,500.00
Common Stock- Dividend         95,000.00
Closing Retained Earnings       364,500.00 AA
Journal Entry
Date Account Debit $ Credit $ Note
2-Jan Cash 6,900,000.00 See E
Common Stock      100,000.00 See F
Paid in capital in excess of par- Common Stock 6,800,000.00 See G
14-Feb Cash      649,000.00 See T
Preferred Stock      590,000.00 See U
Paid in capital in excess of par- Preferred Stock        59,000.00 See V
8-May Treasury Stock      590,000.00
Cash      590,000.00 See M
31-May Cash      320,000.00 See Q
Treasury Stock      295,000.00 See R
Paid in capital in excess of par- Common Stock        25,000.00 See S
1-Dec Retained Earnings      124,500.00
Preferred Stock- Dividend Payable        29,500.00 Given in the sum
Common Stock- Dividend Payable        95,000.00 See J
31-Dec Preferred Stock- Dividend Payable        29,500.00 Given in the sum
Common Stock- Dividend Payable        95,000.00 See J
Cash      124,500.00
31-Dec Income Summary      489,000.00 Given in the sum
Retained Earnings      489,000.00
Stockholder's Equity
Paid-In Capital: Amount $ Note
Preferred Stock (5%, $ 10 par,
59,000 shares issued and outstanding)        590,000.00 See Y
Common Stock ($ 1 par,
100,000 shares issued and outstanding)        100,000.00 See F
Paid in capital in excess of par- Preferred Stock          59,000.00 See Z
Paid-In Capital in Excess of Par-Common Stock     6,825,000.00 See O
Retained Earnings        364,500.00 See AA
Treasury stock
(5,000 shares at $ 59 par share)      (295,000.00) See P
Total Stockholder's Equity    7,643,500.00